OpenStudy (anonymous):

Convert each subtraction problem to addition and type in the matching number. 1. 3/4 - 2/3 = 2. 3/4 - -2/3 = 3. -3/4 - 2/3 = 4. -3/4 - -2/3 = a. -3/4 + -2/3 b. -3/4 + 2/3 c. 3/4 + 2/3 d. 3/4 + -2/3

OpenStudy (kc_kennylau):

just remember: - + => - - - => + + - => - + + => + that'll help you :D

OpenStudy (anonymous):

I'm still a bit confused :/. I'm terribly bad at math.

OpenStudy (kc_kennylau):

trust in yourself :D everybody has potential :D

OpenStudy (anonymous):

Yeah well in this case I could use some help, kinda the reason I posted here. :D I can't figure out how to use what you posted with the problem.

OpenStudy (kc_kennylau):

a+-b=a-b a-+b=a-b a--b=a+b a++b=a+b